# Unveiling Ferroelectric-Like Behavior in Leaky Dielectrics: A Microscopic Model for Polarization Dynamics and Hysteresis Inversion

### v2

@@ III Conclusions
In this work, we have developed a comprehensive model to describe the dielectric response of non-polar, non-ferroelectric materials, with a particular focus on the interplay between polarization transfer and charge leakage mechanisms. By employing a microscopic approach, we derived analytical expressions for the polarization and current dynamics under cyclic voltammetry and impedance spectroscopy, which allow us to probe the complex transport behavior in leaky dielectrics. Our model successfully reproduces key experimental phenomena, including the inversion of polarization loops, as observed in experimental studies, such as those reported in ferroelectric-like materials. This polarization loop inversion, driven by changes in the voltage sweep period and the dominant contribution of either polarization transfer or leakage, underscores the tunability of the dielectric response.
Furthermore, the model demonstrates how ferroelectric-like hysteresis loops can emerge in non-ferroelectric materials purely due to the presence of non-equilibrium dipoles and energy-dependent transfer rates, providing new insights into the mimicry of ferroelectric traces induced by the presence of defects. Our findings closely match experimental observations of clockwise and counterclockwise reversible polarization loops, which are often misinterpreted as evidence of ferroelectricity. By clearly distinguishing these effects from true ferroelectric behavior, our study not only offers a theoretical foundation for such experimental results but also contributes to the accurate characterization of materials used in capacitors, transistors, and memory devices.
The model’s flexibility in incorporating higher-order harmonics and its extension to multimode impedance analysis further enrich the understanding of dielectric systems dominated by displacement currents. Ultimately, this work provides a valuable framework for interpreting complex dielectric responses, paving the way for more precise experimental characterization and improved device design in systems where distinguishing between real and apparent ferroelectric behavior is crucial.
@@ IV Appendix
Acknowledgements.

### v3

@@ III Conclusions
In this work, we developed a comprehensive model to describe the dielectric response of non-polar, non-ferroelectric materials, focusing on the interplay between polarization transfer and charge leakage mechanisms. Our approach successfully reproduces key experimental phenomena, including the inversion of polarization loops and the emergence of ferroelectric-like hysteresis in leaky dielectrics. By differentiating true ferroelectric behavior from effects induced by non-equilibrium dipoles and energy-dependent transfer rates, the model provides a theoretical foundation for experimental observations and informs the accurate characterization of materials used in capacitors, transistors, and memory devices.
In capacitors, the model can guide the design of materials with reduced leakage and enhanced energy storage capabilities. For transistors, it can guide the development of high-performance FETs by optimizing gate dielectrics for improved control and reduced leakage. Furthermore, the model can contribute to the design of more reliable and efficient resistive RAMs and ferroelectric RAMs by providing insights into the factors influencing the stability and reliability of memristive and ferroelectric states.
Notably, the model highlights the tunability of dielectric responses through parameters such as voltage sweep rate and the dominance of polarization or leakage processes, offering insights into device optimization. Its ability to incorporate higher-order harmonics and multimode impedance analysis enhances its relevance for analyzing systems dominated by displacement currents.
Future extensions could explore material-specific adaptations, the effects of temperature on transport dynamics, and the role of external stimuli such as illumination and applied electro-magnetic fields. These advancements would further refine the model’s applicability to diverse dielectric systems and its utility in guiding experimental protocols and the design of electronic devices.
@@ IV Appendix
$V$ Applied voltage bias
$\eta$ Local voltage efficiency
$L$ Length of the charge localization site
$P$ Electric polarization
$\Omega=D\cdot A$ System volume = system length $\cdot$ system area
$\partial q$ Non-equilibrium charge fluctuation
$F^{(p)}$ Non-equilibrium carrier flux contributing to space charge accumulation
$F^{(lk)}$ Non-equilibrium carrier flux contributing to charge leakage
$\alpha$ Symmetry control parameter
$T_{eff}$ Effective temperature
$E^{(p)}$ Effective barrier for carrier activation contributing to space charge accumulation
$E^{(lk)}$ Effective barrier for carrier activation contributing to charge leakage
$\chi$ Electric susceptibility
$E$ Electric field
$\tau$ Relaxation time
$j$ Current density
$I$ Current
$C$ Geometric capacitance
$G$ Conductance
$B$ Susceptance
$Z$ Impedance
$Y$ Admittance
Table 2: List of symbolsn
Acknowledgments